OpenStudy (anonymous):

THe equation of a tangent line to the graph of y=lnx at x=e^2 I get the derivative but then to get the equation I plus it in y-lnx=1/e^2(x-e^2) and that gives me y=x/e^2-1+lnx but my book says it does not have a lnx in the final answer help

OpenStudy (campbell_st):

oops... the point the tangent passes through is \[x = e^2\] then \[y = \ln(e^2)\] you need to simplify this.... the slope of the tangent is correct

OpenStudy (anonymous):

so does that mean I derive ln(e^2)?

OpenStudy (campbell_st):

it means the point you use for the tangent is \[(e^2, \ln(e^2)).... or... (e^2, 2)\]